The speed of light is 3 x 108 meters per second (m/s). What is the speed...

  1. The speed of light is 3 x 108 meters per second (m/s). What is the speed of light in miles per year (mi/yr)?
Answer #1

1 Mile per year is equal to 5.099 808 118 x 10-5 meters per second

now 3 * 10^8 m/s = 3*10^8 * 5.099*10^-5 = 15.297 * 10^3 miles /year